OpenStudy (credmond):

A. Find the critical numbers B. the largest open intervals where the function is increasing C. the largest open intervals where it is decreasing y = x (sq root of 9 - x^2)

OpenStudy (anonymous):

same problem

OpenStudy (anonymous):

\[f(x)=x\sqrt{9-x^2}\] right?

OpenStudy (credmond):

Yeah! I'm having an issue now with after you get from 1/2x (9-x) ^ -1/2

OpenStudy (credmond):

so basically after finding the derivative

OpenStudy (anonymous):

whoa what did you get for the derivative?

OpenStudy (credmond):

so I went from x (9-x) ^1/2

OpenStudy (anonymous):

this is a product, you have to use the product rule \[(fg)'=f'g+g'f\]

OpenStudy (credmond):

ohhhhh

OpenStudy (anonymous):

skip the rational exponent, it is not your friend at all in doing this problem

OpenStudy (credmond):

oh okay!

OpenStudy (anonymous):

you have \[(f(x)=x, f'(x)=1,g(x)=\sqrt{9-x^2},g'(x)=?\] do you know how to find \(g'(x)\) ?

OpenStudy (credmond):

so g'(x) = (9-x^2)^1/2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what you did was to rewrite \(g(x)\) with a rational exponent \[g(x)=\sqrt{9-x^2}=(9-x)^{\frac{1}{2}}\] that is not the derivative it also does not really help you get it

OpenStudy (credmond):

oh okay. can I use that rewritten form to find the derivative though?

OpenStudy (credmond):

oh wait.. nvm

OpenStudy (anonymous):

yes, that and the chain rule but like i said, i wouldn't recommend it

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{x}=\frac{1}{2\sqrt{x}}\] so by the chain rule \[\frac{d}{dx}\sqrt{9-x^2}=\frac{1}{2\sqrt{9-x^2}}\times (-2x)=-\frac{x}{\sqrt{9-x^2}}\]

OpenStudy (anonymous):

if you want to set something equal to zero and solve you have to use radicals, not negative rational exponents, which will only confuse matters

OpenStudy (credmond):

oh okay. gotcha.

OpenStudy (credmond):

so from there do you use the product rule to find the critical numbers?